Please show step by step BA II plus calculator

You buy a 20-year bond with a coupon rate of 7.6%
that has a yield to maturity of 9.5%. (Assume a
face value of $1,000 and semiannual coupon
payments.) Six months later, the yield to maturity
is 10.5%. What is your return over the 6 months?
(Do not round intermediate calculations. Enter
your answer as a percent rounded to 2 decimal
places. Negative amount should be indicated by a
minus sign.)
Answer #1

Calculating Purchase Price of Bond,

Using TVM Calculation,

PV = [FV = 1,000, PMT = 38, N = 40, I = 0.095/2]

PV = $831.25

Calculating Selling Price of Bond,

Using TVM Calculation,

PV = [FV = 1,000, PMT = 38, N = 39, I = 0.105/2]

PV = $761.35

Return Rate = (761.35 - 831.25 + 38)/831.25

Return Rate = -3.84%